QUESTION 9

  1. Jonah has the choice of paying Rita $20,000 today or $80,000 in ten years. Assume Jonah can earn a 10 percent after-tax rate of return. What is the present value of the $80,000 payment due in 10 years?

QUESTION 10

  1. Komiko Tanaka invests $13,000 in LymaBean, Inc. LymaBean does not pay any dividends. Komiko projects that her investment will generate a 11 percent before-tax rate of return. She plans to invest for the long term.    How much cash will Komiko retain, after-taxes, if she holds the investment for 6 years and then she sells it when the long-term capital gains rate is 15 percent?

QUESTION 11

  1. Komiko Tanaka invests $13,000 in LymaBean, Inc. LymaBean does not pay any dividends. Komiko projects that her investment will generate a 11 percent before-tax rate of return. She plans to invest for the long term. If she holds the investment for 6 years and then she sells it when the long-term capital gains rate is 15 percent, what is her after-tax rate of return?
Answer #1

Dear student, only one question is allowed at a time. I am answering the first question

9)

Present value factor

= 1 / ( 1 + Rate of Interest ) ^ Number of years

Where,

Rate of Interest = 10% or 0.10

So, PV Factor for 10 years will be

= 1 / ( 1.10 ) ^ 10

= 1 / 2.593742

= 0.385543

So, Present value

= Amount payable x Present value factor

= $80,000 x 0.385543

= $ 30,843.46